# A quasilinear elliptic equation with absorption term and Hardy potential

### v1

@@ Abstract
$-\Delta_{p}u+\mu\frac{u^{p-1}}{\left|x\right|^{p}}+\left|x\right|^{\theta}u^{q}=0,\qquad x\in\mathbb{R}^{N}\backslash\left\{0\right\}$
@@
MSC 2010: 35J75, 35J92
Keywords: p-Laplacian, Hardy potentials, absorption, singularities, global solutions
@@ 1 Introduction
$-\Delta u+\mu\frac{u}{\left|x\right|^{2}}+b(x)g(u)=0$
@@
twith $\mu<-(\frac{N-2}{2})^{2}$ and $\theta+2>0$, Wei and Du [30] gave a precise behaviour near 00, and a uniqueness result of global solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$. The problem of existence and radiality of global solutions was then solved in [11] for any value of $\theta$, by using thin techniques of supersolutions and subsolutions.
@@
$-\Delta_{p}u+\mu\frac{u^{p-1}}{\left|x\right|^{p}}=u^{q}$
@@
Here our purpose is to extend to equation (1.1) for any $p>1$ and $q>p-1$, the study of [9], [11] relative to the case $p=2$, and simplify the proofs, by a quite different approach. 0ur line of attack is new, even in the case of equation (1.4). It is built on a complete study of the radial case by phase-plane techniques. Indeed, equation (1.1) is invariant by the scaling $T_{k}$ defined for any $k>0$ and $x\in\mathbb{R}^{N}\backslash\left\{0\right\}$ by
@@ 2.2 Local and global behaviour of the solutions
Our results are given following the different assumptions (2.7)
@@
$\lim_{\left|x\right|\rightarrow 0}\left|x\right|^{\gamma}u=a^{\ast},$ (2.9)
@@
$\left\{\begin{array}[]{c}\lim_{r\rightarrow 0}\left|x\right|^{S_{1}}u=k_{1}>0\text{ if }\mu>\mu_{0}\\ \lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\left|x\right|\right|^{\frac{2}{p}})u(x)=l>0,\text{ if }\mu=\mu_{0}\end{array}\right.$ (2.10)
@@
$\lim_{\left|x\right|\rightarrow 0}\left|x\right|^{S_{2}}u=k_{2}>0.$ (2.11)
@@
(iii) There exists global solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$, all of them are radial:
@@
$\bullet$ if $\mu>\mu_{0}$, for any $k_{1}>0$,there exist a unique global solution such that
@@
$\bullet$ if $\mu=\mu_{0}$ there exists a unique global solution such that
$\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{p}}u(x)=l,\qquad\lim_{\left|x\right|\rightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}.$ (2.14)
@@
$\lim_{\left|x\right|\rightarrow 0}\left|x\right|^{\gamma}u=a^{\ast}.$ (2.15)
@@
$\left\{\begin{array}[]{c}\lim_{\left|x\right|\rightarrow\infty}\left|x\right|^{S_{2}}u=k_{2}>0\text{\qquad if }\mu>\mu_{0},\\ \lim_{x\longrightarrow\infty}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\left|x\right|\right|^{\frac{2}{p}})u(x)=\ell>0\qquad\text{ if }\mu=\mu_{0},\end{array}\right.$ (2.17)
@@
$\lim_{x\longrightarrow 0}\left|\ln\left|x\right|\right|^{-\frac{p-1}{q+1-p}}u$ $=$ $\delta_{N,p,q}>0,\qquad\text{if }\gamma=S_{2}=0.$
@@
$\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{S_{1}}(\left|\ln\left|x\right|\right|)^{-\frac{1}{q+1-p}}u$ $=$ $\beta_{N,p,q},\qquad\text{if }\gamma=S_{1}.$
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{q+1-p}}u=c_{N,p,q}>0,\text{ (resp. }\lim_{\left|x\right|\longrightarrow\infty}r\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{q+1-p}}u=c_{N,p,q}).$
@@
$\mathbf{q}_{1}=p-1+\frac{p+\theta}{S_{1}};\qquad\mathbf{q}_{2}=p-1+\frac{p+\theta}{S_{2}},\quad\text{if }\mu\neq 0.$
@@
$\text{(}\mathcal{H}_{1}\text{)}$  $\mu\geq\mu_{0}\qquad\text{and }1<q<\mathbf{q}_{1},$
$\text{(}\mathcal{H}_{2}\text{)}$ $\Longleftrightarrow 0>\mu\geq\mu_{0}\qquad\text{and }\mathbf{q}_{2}<q,$
$\text{(}\mathcal{H}_{3}\text{)}$  $(\mu\geq 0\text{ and }\mathbf{q}_{1}\leq q)\qquad\text{or }(0>\mu>\mu_{0}\text{ and }\mathbf{q}_{1}\leq q\leq\mathbf{q}_{2}),$
$\text{(}\mathcal{H}_{4}\text{)}$ $\Longleftrightarrow\mu=\mu_{0}\qquad\text{and }q=\mathbf{q}_{s};$
@@
$\text{(}\mathcal{H}_{2}\text{)}$$\Longleftrightarrow$ $\mu\leq 0\qquad\text{or (}\mu>0\text{ and }1<q\leq\mathbf{q}_{2}\text{)},$
$\text{(}\mathcal{H}_{3}\text{)}$$\Longleftrightarrow$ $\mu>0\qquad\text{and }\mathbf{q}_{2}\leq q;$
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{\gamma}u=a^{\ast}.\qquad\lim_{\left|x\right|\rightarrow\infty}u=k_{1}>0.$
@@
Moreover, when $\theta=-\frac{p(N-1)}{p-1}$, thus $\theta+p<0$, there exist explicit global and local solutions, in $B_{r_{0}}\backslash\left\{0\right\}$ and in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ of equation (2.23), given by
$u(r)=(c\pm d_{p,q,N}r^{\frac{p-N}{p-1}})^{-\frac{p}{q+1-p}},$
where $c>0$ and $d_{p,q,N}=\frac{(p-1)(q-p+1)}{p(N-p)}(\frac{p}{(p-1)(q+1)})^{\frac{1}{p}}$, see [3, Theorem 4.17].
@@ 4.2 New formulation as an autonomous system
$-(\left|U_{t}\right|^{p-2}U_{t})_{t}+(p-N)\left|U_{t}\right|^{p-2}U_{t}+\mu U^{p-1}+r^{\theta+p}U^{q}=0,$
@@
$\lim_{r\longrightarrow 0}r^{\frac{N-p}{p}}\left|\ln r\right|^{-\frac{2}{p}}u(r)=\ell\qquad\text{(resp. }\lim_{r\longrightarrow\infty}r^{\frac{N-p}{p}}\left|\ln r\right|^{-\frac{2}{p}}u(r)=\ell\text{)}$ (4.11)
@@ 4.3 Main radial results
$\bullet$ There exists a global particular solution $u^{\ast}=a^{\ast}r^{-\gamma}.$
$\bullet$ If $\mu>\mu_{0}$, for any $k_{1}>0$, there exist a unique global solution such that
@@
$\lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{\frac{2}{p}})u(r)=\ell,\qquad\lim_{r\rightarrow\infty}r^{\gamma}u=a^{\ast},\qquad\text{if }\mu=\mu_{0}.$ (4.14)
@@
$\left\{\begin{array}[]{c}\lim_{r\rightarrow 0}r^{S_{1}}u=k_{1}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,\quad\quad\lim_{r\rightarrow R}u=\infty,$ (4.17)
$\left\{\begin{array}[]{c}\lim_{r\rightarrow 0}r^{S_{1}}u=k_{1}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.\text{\qquad}\lim_{r\rightarrow R}u=0,$ (4.18)
@@
$\bullet$ There exists a global particular solution $u^{\ast}=a^{\ast}r^{-\gamma}.$
$\bullet$ If $\mu>\mu_{0}$, for any $k_{2}>0$, there exist a unique global solution such that
@@
 $\lim_{r\rightarrow R}u=\infty,\qquad\left\{\begin{array}[]{c}\lim_{r\rightarrow\infty}r^{S_{2}}u=k_{2}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow\infty}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,$ (4.25)
$\lim_{r\rightarrow R}u=0,\qquad\left\{\begin{array}[]{c}\lim_{r\rightarrow\infty}r^{S_{2}}u=k_{2}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow\infty}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,$ (4.26)
@@
Case ($\mathcal{H}_{4}$) Suppose $\mu=\mu_{0}$ and $\gamma=\frac{N-p}{p}.$ Then there is no global solution. For any $R>0$,there exists solutions such that
@@
Case ($\mathcal{H}_{5}$). Suppose $\mu<\mu_{0}.$ Then there exists a unique global solution: $u=u^{\ast}.$
@@ 5.3 A priori Osserman’s estimate near 00 or $\infty$
$a=\left|x_{0}\right|^{\theta}\min\left\{2^{-(\theta+1)},3.2^{\theta-1}\right\},\qquad b=\left|x_{0}\right|^{\frac{pq+\theta(p-1)}{q-p+1}}\max\left\{2^{-\frac{pq+\theta(p-1)}{q-p+1}},(\frac{3}{2})^{\frac{pq+\theta(p-1)}{q-p+1}}\right\},$
@@ 5.4 Regularity result
Let $q>p-1.$ Let $u$ be a solution of (1.1) in $B_{r_{0}}\backslash\left\{0\right\}$, such that
$u(x)\leq C_{1}\left|x\right|^{-\delta}\qquad\text{in }B_{r_{0}}\backslash\left\{0\right\}$
for some $\delta\leq\gamma$. Then
@@
and there exists $\alpha\in(0,1)$ only depending of $N,p,q,\delta$ such that, for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}$,
@@
Analogous results hold in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ when $\delta\geq\gamma$.
Proof. For $x\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}$ there exists $R\in(0,\frac{r_{0}}{6})$ such that $2R\leq\left|x\right|\leq 3R$. Let $\Gamma=B_{7}\backslash\overline{B_{1}}$ and $\Gamma^{\prime}=B_{6}\backslash\overline{B_{2}}$. We set $u_{R}(\xi)=R^{\delta}u(R\xi)$ for any $\xi\in\Gamma$. Then
$-\Delta_{p}u_{R}(\xi)+\frac{\mu u_{R}^{p-1}(\xi)}{\left|\xi\right|^{p}}+R^{(q+1-p)(\gamma-\delta)}u_{R}^{q}=0$
for $\xi\in\Gamma$. Now $R^{(q+1-p)(\gamma-\delta)}$ stays bounded as $R\rightarrow 0$, thanks to $\delta\leq\gamma$ and $q>p-1;$ and $\left\|u_{R}\right\|_{L^{\infty}(\Gamma)}\leq C_{1}\left|\xi\right|^{\delta}\leq C_{1}^{\prime},$ thus $\Delta_{p}u_{R}$ is bounded in $\Gamma$. Then from Tolksdorf results of [25, Theorem 1], $\left|\nabla u_{R}\right|$ is bounded in $C^{0,\alpha}(\Gamma^{\prime})$ independently of $R$ for some $\alpha\in(0,1)$. Then (5.3) holds. Moreover, let $x,x^{\prime}\in B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}$. First assume $\left|x\right|\leq\left|x^{\prime}\right|\leq 2\left|x\right|$. Then $\xi=\frac{x}{R}$ and $\xi^{\prime}=\frac{x^{\prime}}{R}\in\Gamma^{\prime}$, and
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|=R^{-1-\delta}\left|\nabla u_{R}(\xi)-\nabla u(\xi^{\prime})\right|\leq CR^{-1-\delta}\left|\xi-\xi^{\prime}\right|^{\alpha}\leq C_{2}\left|x\right|^{-(\delta+1+\alpha)}\left|x-x^{\prime}\right|^{\alpha}.$
@@
implying (5.4).
Remark 5.7
In the same way, if $u$ is a solution in $B_{r_{0}}\backslash\left\{0\right\}$, with $r_{0}<1,$ such that
$u(x)\leq C_{1}\left|x\right|^{-\delta}\left|\ln\left|x\right|\right|\qquad\text{in }B_{r_{0}}\backslash\left\{0\right\}$
for some $\delta<\gamma.$ Then for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\},$
$\left|\nabla u(x)\right|\leq C_{2}\left|x\right|^{-(\delta+1)}\left|\ln\left|x\right|\right|$
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|\leq C_{3}(\left|x^{\prime}\right|^{-\delta}+\left|x\right|^{-\delta})\left|x\right|^{-(1+\alpha)}\left|\ln\left|x\right|\right|\left|x-x^{\prime}\right|^{\alpha}.$

### v2

@@ Abstract
$-\Delta_{p}u+\mu\frac{u^{p-1}}{\left|x\right|^{p}}+\left|x\right|^{\theta}u^{q}=0,\qquad x\in\mathbb{R}^{N}\backslash\left\{0\right\},$
@@ 1 Introduction
$-\Delta u+\mu\frac{u}{\left|x\right|^{2}}+b(x)g(u)=0,$
@@
with $\mu<-(\frac{N-2}{2})^{2}$ and $\theta+2>0$, Wei and Du [30] gave a precise behaviour near 00, and a uniqueness result of global solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$. The problem of existence and radiality of global solutions was then solved in [11] for any value of $\theta$, by using thin techniques of supersolutions and subsolutions.
@@
$-\Delta_{p}u+\mu\frac{u^{p-1}}{\left|x\right|^{p}}=u^{q},$
@@
Here our purpose is to extend to equation (1.1) for any $p>1$ and $q>p-1$, the study of [9], [11] relative to the case $p=2$, and simplify the proofs, by a quite different approach. Our line of attack is new, even in the case of equation (1.4). It is built on a complete study of the radial case by phase-plane techniques. Indeed, equation (1.1) is invariant by the scaling $T_{k}$ defined for any $k>0$ and $x\in\mathbb{R}^{N}\backslash\left\{0\right\}$ by
@@ 2.2 Local and global behaviour of the solutions
Our results are given following the different assumptions (2.7):
@@
$\lim_{x\rightarrow 0}\left|x\right|^{\gamma}u=a^{\ast},$ (2.9)
@@
$\left\{\begin{array}[]{ccc}\lim_{x\rightarrow 0}\left|x\right|^{S_{1}}u=k_{1}>0,&&\text{if }\mu>\mu_{0},\\ \lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\left|x\right|\right|^{-\frac{2}{p}})u=\ell>0,&&\text{if }\mu=\mu_{0},\end{array}\right.$ (2.10)
@@
$\lim_{x\rightarrow 0}\left|x\right|^{S_{2}}u=k_{2}>0.$ (2.11)
@@
(iii) There exist global solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$, all of them are radial:
@@
$\bullet$ if $\mu>\mu_{0},$ for any $k_{1}>0$, there exists a unique global solution such that
@@
$\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{-\frac{2}{p}}u=\ell,\qquad\lim_{\left|x\right|\rightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}.$ (2.14)
@@
$\lim_{x\rightarrow 0}\left|x\right|^{\gamma}u=a^{\ast}.$ (2.15)
@@
$\left\{\begin{array}[]{ccc}\lim_{\left|x\right|\rightarrow\infty}\left|x\right|^{S_{2}}u=k_{2}>0,&&\text{if }\mu>\mu_{0},\\ \lim_{\left|x\right|\rightarrow\infty}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{-\frac{2}{p}}u=\ell>0,&&\text{if }\mu=\mu_{0},\end{array}\right.$ (2.17)
@@
$\lim_{x\longrightarrow 0}\left|\ln\left|x\right|\right|^{-\frac{p-1}{q+1-p}}u$ $=$ $\delta_{N,p,q}>0\qquad\text{if }\gamma=S_{2}=0.$
@@
$\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{S_{1}}(\left|\ln\left|x\right|\right|)^{-\frac{1}{q+1-p}}u$ $=$ $\beta_{N,p,q}\qquad\text{if }\gamma=S_{1}.$
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{q+1-p}}u=c_{N,p,q}>0,\text{ (resp. }\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{q+1-p}}u=c_{N,p,q}).$
@@
$\mathbf{q}_{1}=p-1+\frac{p+\theta}{S_{1}},\qquad\mathbf{q}_{2}=p-1+\frac{p+\theta}{S_{2}},\quad\text{if }\mu\neq 0.$
@@
$\text{(}\mathcal{H}_{1}\text{)}$  $\mu\geq\mu_{0}\qquad\text{and\quad\ }1<q<\mathbf{q}_{1},$
$\text{(}\mathcal{H}_{2}\text{)}$ $\Longleftrightarrow 0>\mu\geq\mu_{0}\qquad\text{and\quad\ }\mathbf{q}_{2}<q,$
$\text{(}\mathcal{H}_{3}\text{)}$  $(\mu\geq 0\text{ and }\mathbf{q}_{1}\leq q)\quad\text{or\quad}(0>\mu>\mu_{0}\text{ and }\mathbf{q}_{1}\leq q\leq\mathbf{q}_{2}),$
$\text{(}\mathcal{H}_{4}\text{)}$ $\Longleftrightarrow\mu=\mu_{0}\qquad\text{and}\quad q=\mathbf{q}_{s};$
@@
$\text{(}\mathcal{H}_{2}\text{)}$$\Longleftrightarrow$ $\mu\leq 0\qquad\text{or\quad(}\mu>0\text{ and }1<q\leq\mathbf{q}_{2}\text{)},$
$\text{(}\mathcal{H}_{3}\text{)}$$\Longleftrightarrow$ $\mu>0\qquad\text{and\quad}\mathbf{q}_{2}\leq q;$
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{\gamma}u=a^{\ast},\qquad\lim_{\left|x\right|\rightarrow\infty}u=k_{1}>0.$
@@
Moreover, when $\theta=-\frac{p(N-1)}{p-1}$, thus $\theta+p<0$, there exist explicit of equation (2.23), given by
$\0u(r)=(c+\epsilon d_{p,q,N}r^{\frac{p-N}{p-1}})^{-\frac{p}{q+1-p}},\qquad\epsilon=\pm 1,\quad c\in\mathbb{R}\backslash\left\{0\right\},$
where $d_{p,q,N}=\frac{(p-1)(q-p+1)}{p(N-p)}(\frac{p}{(p-1)(q+1)})^{\frac{1}{p}},$ global if $c>0$ and $\epsilon=1,$ local near 00 if $c<0$ and $\epsilon=1,$ or near $\infty$ if $c>0$ and $\epsilon=-1,$ see also [3, Theorem 4.17].
@@ 4.2 New formulation as an autonomous system
$-(\left|U_{t}\right|^{p-2}U_{t})_{t}+(p-N)\left|U_{t}\right|^{p-2}U_{t}+\mu U^{p-1}+r^{\theta+p}U^{q}=0.$
@@
$\lim_{r\longrightarrow 0}r^{\frac{N-p}{p}}\left|\ln r\right|^{-\frac{2}{p}}u(r)=\ell\qquad(\text{resp. }\lim_{r\longrightarrow\infty}r^{\frac{N-p}{p}}\left|\ln r\right|^{-\frac{2}{p}}u(r)=\ell).$ (4.11)
@@ 4.3 Main radial results
$\bullet$ There exists a global particular solution $u^{\ast}=a^{\ast}r^{-\gamma}$ of (4.1).
$\bullet$ If $\mu>\mu_{0}$, for any $k_{1}>0$, there exist a unique global solution of (4.1) such that
@@
$\lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{-\frac{2}{p}})u(r)=\ell,\qquad\lim_{r\rightarrow\infty}r^{\gamma}u=a^{\ast},\qquad\text{if }\mu=\mu_{0}.$ (4.14)
@@
$\left\{\begin{array}[]{c}\lim_{r\rightarrow 0}r^{S_{1}}u=k_{1}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{-\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,\quad\quad\lim_{r\rightarrow R}u=\infty,$ (4.17)
$\left\{\begin{array}[]{c}\lim_{r\rightarrow 0}r^{S_{1}}u=k_{1}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow 0}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{-\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.\text{\qquad}\lim_{r\rightarrow R}u=0,$ (4.18)
@@
$\bullet$ There exists a global particular solution $u^{\ast}=a^{\ast}r^{-\gamma}$ of (4.1).
$\bullet$ If $\mu>\mu_{0}$, for any $k_{2}>0$, there exist a unique global solution of (4.1)such that
@@
 $\lim_{r\rightarrow R}u=\infty,\qquad\left\{\begin{array}[]{c}\lim_{r\rightarrow\infty}r^{S_{2}}u=k_{2}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow\infty}r^{\frac{N-p}{p}}(\left|\ln(r)\right|-^{\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,$ (4.25)
$\lim_{r\rightarrow R}u=0,\qquad\left\{\begin{array}[]{c}\lim_{r\rightarrow\infty}r^{S_{2}}u=k_{2}>0\text{ if }\mu>\mu_{0}\\ \lim_{r\rightarrow\infty}r^{\frac{N-p}{p}}(\left|\ln(r)\right|^{-\frac{2}{p}})u(r)=\ell>0,\text{ if }\mu=\mu_{0}\end{array}\right.,$ (4.26)
@@
Case ($\mathcal{H}_{4}$) Suppose $\mu=\mu_{0}$ and $\gamma=\frac{N-p}{p}.$ Then there is no global solution of (4.1). For any $R>0$,there exists solutions such that
@@
Case ($\mathcal{H}_{5}$). Suppose $\mu<\mu_{0}.$ Then there exists a unique global solution: $u^{\ast}=a^{\ast}r^{-\gamma}$ of (4.1).
@@ 5.3 A priori Osserman’s estimate near 00 or $\infty$
$a=2^{-(\theta+1)}\min\left\{1,3^{\theta}\right\}\left|x_{0}\right|^{\theta}\qquad b=\left|x_{0}\right|^{\frac{pq+\theta(p-1)}{q-p+1}}\max\left\{2^{-\frac{pq+\theta(p-1)}{q-p+1}},(\frac{3}{2})^{\frac{pq+\theta(p-1)}{q-p+1}}\right\},$
@@ 5.4 Regularity result
Let $q>p-1.$ Let $u$ be a solution of (1.1) in $B_{r_{0}}\backslash\left\{0\right\}$(resp. $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}})$ such that
$u(x)\leq C_{1}\left|x\right|^{-\delta}\qquad\text{in }B_{r_{0}}\backslash\left\{0\right\}\quad(\text{resp.}\mathbb{R}^{N}\backslash\overline{B_{r_{0}}})$
for some $\delta\leq\gamma$ (resp. $\delta\geq\gamma).$ Then
@@
and there exists $\alpha\in(0,1)$ only depending of $N,p,q,\delta$ such that, for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}$ ((((resp. $\mathbb{R}^{N}\backslash\overline{B_{2r_{0}}})$
@@
Proof. Let $u$ be a solution of (1.1) in $B_{r_{0}}\backslash\left\{0\right\}.$ By the scaling (1.7) we can assume $r_{0}=1.$ Let $\Gamma=B_{7}\backslash\overline{B_{1}}$ and $\Gamma^{\prime}=B_{6}\backslash\overline{B_{2}}$. For any $R\in(0,\frac{1}{7}),$ we set $u_{R}(\xi)=R^{\delta}u(R\xi)$ for any $\xi\in\Gamma$. Then
$-\Delta_{p}u_{R}(\xi)+\frac{\mu u_{R}^{p-1}(\xi)}{\left|\xi\right|^{p}}+R^{(q+1-p)(\gamma-\delta)}u_{R}^{q}(\xi)=0$
for $\xi\in\Gamma$. Then $u_{R}$ is bounded in $\Gamma:$
$u_{R}(\xi)\leq C_{1}\left|\xi\right|^{\delta}\leq C_{1}\max\left\{1,7^{\delta}\right\}.$
Moreover $R^{(q+1-p)(\gamma-\delta)}$ stays bounded on $(0,\frac{1}{7})$, thanks to the assumptions on $\delta$ and $q,$ thus $\Delta_{p}u_{R}$ is bounded in $\Gamma$. Then from Tolksdorf results of [25, Theorem 1], $\left|\nabla u_{R}\right|$ is bounded in $C^{0,\alpha}(\Gamma^{\prime})$ independently of $R$ for some $\alpha\in(0,1)$. For $x\in$ $B_{\frac{1}{2}}\backslash\left\{0\right\}$ there exists $R\in(0,\frac{1}{7})$ such that $2R<\left|x\right|<6R$.Then (5.3) holds. Moreover, let $x,x^{\prime}\in B_{\frac{1}{2}}\backslash\left\{0\right\}$. First assume $\left|x\right|\leq\left|x^{\prime}\right|\leq 2\left|x\right|$. Then $\xi=\frac{x}{R}$ and $\xi^{\prime}=\frac{x^{\prime}}{R}\in\Gamma^{\prime}$, and
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|=R^{-1-\delta}\left|\nabla u_{R}(\xi)-\nabla u_{R}(\xi^{\prime})\right|\leq CR^{-1-\delta}\left|\xi-\xi^{\prime}\right|^{\alpha}\leq C_{2}\left|x\right|^{-(\delta+1+\alpha)}\left|x-x^{\prime}\right|^{\alpha}.$
@@
implying (5.4). We get analogous results hold in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ when $\delta\geq\gamma.\vskip 6.0pt plus 2.0pt minus 2.0pt$
Moreover in the case of a logarithmic type estimate, we obtain the following:
Proposition 5.7
Let $q>p-1$ and $s>0.$ Let $u$ be a solution of (1.1) in $B_{r_{0}}\backslash\left\{0\right\}$ with $r_{0}\leq 1,$ (resp. in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ with $r_{0}\geq 1$) such that
$u(x)\leq C_{1}\left|x\right|^{-\delta}(\left|\ln\left|x\right|\right|+1)^{s}\qquad\text{in }B_{r_{0}}\backslash\left\{0\right\}$
for some $\delta<\gamma$ (resp. $\delta>\gamma).$ Then for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}$ (resp. $\mathbb{R}^{N}\backslash\overline{B_{2r_{0}}}$),
$\left|\nabla u(x)\right|\leq C_{2}\left|x\right|^{-(\delta+1)}(\left|\ln\left|x\right|\right|+1)^{s}$
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|\leq C_{3}(\left|x^{\prime}\right|^{-\delta}+\left|x\right|^{-\delta})\left|x\right|^{-(1+\alpha)}(\left|\ln\left|x\right|\right|+1)^{s}\left|x-x^{\prime}\right|^{\alpha}.$
Proof. As above we can assume $r_{0}=1.$ Let $u$ be a solution of (1.1) in $B_{1}\backslash\left\{0\right\}.$Taking $u_{R}(\xi)=R^{\delta}\left|\ln R\right|^{-s}u(R\xi)$ for any $\xi\in\Gamma,$ we obtain
$-\Delta_{p}u_{R}(\xi)+\frac{\mu u_{R}^{p-1}(\xi)}{\left|\xi\right|^{p}}+R^{(q+1-p)(\gamma-\delta)}\left|\ln R\right|^{s(q-p+1)}u_{R}^{q}(\xi)=0.$
The function $u_{R}$ is still bounded in $\Gamma$ because $s\geq 0:$
$u_{R}(\xi)\leq C_{1}\left|\xi\right|^{\delta}(\frac{\left|\ln(R\xi)\right|+1}{\left|\ln R\right|})^{s}\leq C_{1}\left|\xi\right|^{\delta}(1+\frac{\ln\xi+1}{\left|\ln R\right|})^{s}\leq C_{1}\max\left\{1,7^{\delta}\right\}(1+\frac{\ln 7+1}{\ln 7})^{s}.$
And $R^{(q+1-p)(\gamma-\delta)}\left|\ln R\right|^{s(q-p+1)}$ is bounded on $(0,\frac{1}{7}),$ because $\delta<\gamma.$ Then we conclude as before, and similarly in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}.$